# Learning Dynamic Bayesian Networks from Data: Foundations, First Principles and Numerical Comparisons

### v1

@@ 2.1 Simple Parametric Conditional Probability Dependency
For certain kinds of variables, it becomes both possible and prudent to use certain simple parametric families for defining $f$. In particular, for binary Bernoulli random variables correspond to Dirichlet distributions for the prior of the weights together with using Conditional Probability Tables to define the form. For continuous random variables, Gaussian linear models provide a means of computing the maximum likelihood linear parameters using covariance matrices.
A significant advantage of using parametric families arises from the closed form computation of criteria, which permits closed form computation of the marginal posterior of a structure. This permits structure learning algorithms to be able to score graphs offline, assisting in the search. Many score maximizing procedures such as [26, 14, 3, 2] use this approach. The score is ultimately an integration of the posterior of the parameters in the model given the structure, which also indicates that the sampling of the optimal parameters, once obtaining the maximum a posteriori structure, is straightforward for these models.
In addition, one can use neural models including the Generative Flow Network approaches as given by [17, 2]. These use a Reinforcement Learning iteration to ultimately sample from a high scoring network as according to a defined score. Reinforcement Learning broadly, e.g., [74] is another framework by which the structure search for these standard specific models can be aided by neural networks.
@@ 2.1.1 Discrete Variables
The case of binary random variables is wherein $X_{i}(t)\in B(1,p^{X}_{i}),Z_{j}\sim B(1,p^{Z}_{j})$, etc., that is, they are all of Bernoulli type. Empirical samples for all $k\in[K]$, where $k$ indexes a set of sample trajectories satisfy $X^{(k)}_{i}(t),Z^{(k)}_{j}\in\{0,1\}$ for all $t$, $i$, and $j$. With this most simple scenario, the modeling flexibility as well as the nuances of structure learning becomes a natural pedagogical start.
There is a degree of flexibility in the choice of statistical model for defining the transition function. We will explore three options - the noisy or model, the linear logit model, and the complete linear logit model.
@@
$X_{i}(t+1)=\sum\limits_{j\in dpa_{d}(i)}\beta^{d}_{j}X_{j}(t)+\sum\limits_{j\in dpa_{s}(i)}\beta^{s}_{j}X_{j}(t+1)+\sum\limits_{\tau\in dpa_{\tau}(i)}\beta^{a}_{\tau}X_{i}(t-\tau)+\sum\limits_{j\in dpa_{z}(i)}\beta^{z}_{j}Z_{j}$ (5)
@@
$p(X_{i}(t+1)=1)=\sigma\left(\beta_{0}+\sum\limits_{j\in dpa_{d}(i)}\beta^{d}_{j}X_{j}(t)+\sum\limits_{j\in dpa_{s}(i)}\beta^{s}_{j}X(t+1)_{j}+\sum\limits_{\tau\in dpa_{\tau}(i)}\beta^{a}_{\tau}X_{i}(t-\tau)+\sum\limits_{j\in dpa_{z}(i)}\beta^{z}_{j}Z_{j}\right)$ (6)
@@
$\begin{array}[]{l}X_{1}(t+1)=\beta_{0}+\beta_{1}Z_{2}+\beta_{2}X_{3}(t+1)+\beta_{3}X_{3}(t+1)Z_{2}+\beta_{4}X_{3}(t+1)Z_{2}+\beta_{5}X_{2}(t)\\ \qquad\qquad\qquad+\beta_{6}X_{2}(t)Z_{2}+\beta_{7}X_{2}(t)X_{3}(t+1)+\beta_{8}X_{2}(t)X_{3}(t+1)Z_{2}\end{array}$
@@
$\begin{array}[]{l}X_{i}(t+1)=\prod\limits_{j\in dpa_{d}(i)}\prod\limits_{k\in dpa_{s}(i)}\prod\limits_{\tau=1,...,p}\prod\limits_{l\in dpa_{z}(i)}\sum_{\alpha\in\left(\mathbb{Z}^{+}_{2}\right)^{4}}\beta_{i,j,k,l,\tau}^{\alpha}\left(X_{j}(t)X_{k}(t+1)X_{i}(t-\tau)Z_{j}\right)^{\alpha}\end{array}$ (7)
@@
On the other hand, this structure of statistical model presents two structural conditions denoted as local parameter independence and unrestricted multinomial distribution [61]. These ensure that for every configuration, that is, every possible combinations of values of the parents of a node, there is an independent parameter vector. This leads to a corresponding combinatorial explosion of parameter vectors in the statistical model. On the other hand, however, the analytical calculation of parameter likelihoods and posterior distributions become possible, facilitating more straightforward evaluation of scoring metrics quantifying the information quality of an entire (D)BN. That is, the marginal likelihood of a structure can be computed without first computing the likelihood of the weights.
@@
$\sum\limits_{j\in dpa_{d}(i)}\beta^{d}_{j}X_{j}(t)+\sum\limits_{j\in dpa_{s}(i)}\beta^{s}_{j}X_{j}(t+1)+\sum\limits_{\tau\in dpa_{\tau}(i)}\beta^{a}_{\tau}X_{i}(t-\tau)+\sum\limits_{j\in dpa_{z}(i)}\beta^{z}_{j}Z_{j}=\sum_{j\in dpa(i)}\beta_{j}V_{j}(t+1)$
@@
$p(X_{i}(t+1)=u^{l}|V_{dpa(i)}(t+1),\theta_{i}):=\theta_{i}^{\xi(V_{dpa(i)}(t+1)},\,\xi\in\Xi_{i},\,\Xi_{i}=\prod_{j\in dpa_{t}(i)}\mathbb{Z}_{m}^{+}\times\prod_{j\in dpa_{z}(i)}\mathbb{Z}_{q}^{+}$ (9)
That is, there is an multi-index that enumerates the entries of $\Xi_{i}$ for each transition $i$. We can see that this presents a highly parametrized model, which will imply significant parametric uncertainty when there are finite data samples. On the other hand, with this highly precise model, the maximum likelihood becomes much more straightforward to compute, as well as the Bayesian scores. Indeed, this is exactly what local parameter independence facilitates – you can compute the likelihood by counting the instances of each transition and dividing by the count of each predecessor configuration. On the other hand, when the total count of every possible predecessor configuration is low, due to unfavorable sample complexity, this cannot be said to be a high quality estimate of the actual validity of that dependence. On the other hand, by letting these parameters take on distributions, in a Bayesian setting, the computation of a posterior for a structure becomes easier, and the uncertainty is available by sampling the posterior, anyway.
@@ 2.1.2 Continuous Variables
Indeed in [36, Theorem 7.3-7.4] it is shown that there is a bidirectional equivalence between such a normal joint distribution and normal transition function.
@@
Linear Structural Equation Models (LSEMs) are the most common non-Gaussian DBN for modeling continuous variables. With LSEMs (see, e.g. [7] for a general reference and [53] for application to causal inference) presume a general linear structure that is associated with a discretization of a dynamical system:
@@
An SEM could refer to a purely time-instant (static) model, with dependencies $dpa_{s}$ and $dpa_{z}$ only, as in [17, 45]. More recently, DBNs more broadly have become interchangeable with SEMs, for instance, the representation in [50] has all dependencies as described here except, it can be argued for simplicity, $Z$.
@@
Broadly speaking, there are a number of popular parametric forms of nonlinear models that can be used from time series literature, e.g. [19]. Neural networks have enabled computationally intensive empirical unsupervised time series models [22, 9]. Nonlinear models in the SEM statistical community have also been studied [42]. The work [58] uses splines to model the nonlinear relationships in the transition distributions. The work [35] uses a kernel nonparametric regression model to learn DBNs for gene regulatory networks.
@@ 3 Learning From Multiple Trajectories
and we are interested in fitting a DBN model to this data. This amounts to defining the specific form of $f$ in (1). More specifically, it amounts to identifying the parents of each $X_{j}(t)$ in the graph $\bar{\mathcal{G}}$, as well as specific functional form of the transition function $f$.
@@ 3.1 Maximum Likelihood Calculations
In reviewing the literature on learning DBNs from data, it is typical to disregard the distinction of the trajectory sample $\mathcal{T}^{i}$ and the time transition samples $\{X^{(i)}(t),X^{(i)}(t+1)\}$. As far as understanding the meta-methodological cause of this, it appears that this can be said to be due to DBNs being considered not uniquely, but as a special kind of Bayesian Network, or as splices of the same time series trajectory.
Consider the two methodological components thereof, time series analysis and PGMs. For the latter, consider two popular works that are effectively extensions of methods developed for BNs extended to DBNs, the continuous reformulation of the problem into one with adjacency matrices as decision variables, called “NOTEARS” in the static case [72] and “dynotears” [50], as well as the use of “Generative Flow Networks”, a Reinforcement Learning-motivated sampler, for the static case in [17] and the dynamic case in [2]. It can be seen that in all of these cases, the likelihood is expressed as,
@@
And similarly, in consulting standard texts on time series analysis with detailed derivations of Likelihood computation for various models, e.g. [19, 47], we see that in the derivations of the likelihood, the data is considered to be a sequence of i.i.d. observations, that is, a sequence of observations from a stochastic process $\{\hat{X}(0),\hat{X}(1),\hat{X}(2),...,\hat{X}(T)\}$, rather than the general form given in (12), and is fit to (13), just with a simpler expression in the sum index.
This presents the natural question as to whether or, in light of this expression’s universal use, why, these approaches “commute”, that is, whether the equivalent expressions for the likelihood, brought from different perspectives, are appropriately equivalent and true.
We shall see that indeed, arithmetically, the expression for the likelihood is correct for DBNs, and so this makes the calculation of the
3.2 Considerations from Axioms of Causal Learning
DBNs, compared to BNs, contain both time-varying as well as static variables. While the aggregate structure is still a DAG, suggesting that formally many of the same principles regarding inference as well as structure and weight learning in BNs carry over to DBNs, the presence of time, especially when long trajectories are expected, adds significant complications.
This indicates the utility of including both static $Z$ and dynamic $X$ variables in the model. Instead one can consider a new trajectory as a new sample of $\hat{Z}$, which itself samples $X(0)\sim\pi(X(0)|Z)$ then, $X(0),X(1),...,X(T)$. As such, one has $T$ samples in order to learn $P\left(X(t+1)|X(t),\hat{Z}\right)$. However, what can be said about how informative a marginal trajectory is towards learning $P\left(X(t+1)|X(t)\right)$, that is, the marginal conditional over the population of $Z$?
As far the required observations for causal identification, there exists at least one $Z$ such that for all trajectories $Z$ is observed, and $Z$ is in the parent of some $X(t)$. We can consider that the classic Randomized Clinical Trial is exactly that $Z_{H}\in\{0,A\}$ and then testing for $p(X(t+1)|X(t),Z_{\setminus},Z_{H}=0)\neq p(X(t+1)|X(t),Z_{\setminus},Z_{H}=A)$, with a null and alternative hypothesis and $Z_{\setminus}$ as other covariates, assumed to me completely independent of $Z_{H}$.
3.2.1 Conditional Independence and d(irected)-separation
@@
Independence of various sets of variables can be determined by examining d-separation (d means directional) criterion of the (D)BN dag [13].
@@
3.2.2 Causal Sufficiency
We discuss various definition from [5].
@@ 3.2.2 Causal Sufficiency
3.2.3 Causal discovery and Inference
@@ 3.3 Closed System Graph Causal Identification Model and Likelihood Information
wherein $\epsilon$ is a zero mean error term. This additive noise model formulation has been leveraged to establish results on the identifiability of the structure $\bar{\mathcal{G}}$ [53, 32].
@@
Finally, let us investigate in more detail the validity of the iid-inter-intra-trajectory assumption implicit in the likelihood form given by (13).
@@
from which we can see the intuition of the circumstance. Asymptotically, the second term approaches the stationary distribution, and the independence assumption becomes valid. Otherwise, we can consider that for $T$ much longer than the mixing time, this assumption is also valid for most of the transitions. However, otherwise we can see that:
@@
    where $V_{i}$ could be any predecessor in the graph for $X_{i}(t+1)$.

### v2

@@ 2.1 Simple Parametric Conditional Probability Dependency
For certain kinds of variables, it becomes both possible and prudent to use certain simple parametric families for defining $f$. For instance, for binary Bernoulli random variables, the use of Dirichlet distributions for the prior of the weights together with using Conditional Probability Tables to define the form. For continuous random variables, Gaussian linear models provide a means of computing the maximum likelihood linear parameters using covariance matrices.
A significant advantage of using parametric families arises from the closed form computation of criteria, which permits closed form computation of the marginal posterior of a structure. This permits structure learning algorithms to be able to score graphs offline, assisting in the search. Many score maximizing procedures such as [geiger2002parameter, cussens2020gobnilp, BARTLETT2017258, atanackovic2024dyngfn] use this approach. The score is ultimately an integration of the posterior of the parameters in the model given the structure, which also indicates that the sampling of the optimal parameters, once obtaining the maximum a posteriori structure, is straightforward for these models.
With the optimal parameters and likelihood computation amortized, one can then use black box approaches such as those involving neural models. This includes the innovative Generative Flow Network approaches as given by [pmlr-v180-deleu22a, atanackovic2024dyngfn]. These use a Reinforcement Learning iteration to ultimately sample from a high scoring network as according to a defined score. Reinforcement Learning broadly, e.g., [zheng2023rbnets] is another framework by which the structure search for these standard specific models can be aided by neural networks.
@@ 2.1.1 Discrete Variables
The case of binary random variables is wherein $X_{i}(t)\in B(1,p^{X}_{i}(t)),Z_{j}\sim B(1,p^{Z}_{j})$, etc., that is, they are all of Bernoulli type. Empirical samples for all $k\in[K]$, where $k$ indexes a set of sample trajectories satisfy $X^{(k)}_{i}(t),Z^{(k)}_{j}\in\{0,1\}$ for all $t$, $i$, and $j$. With this most simple scenario, the modeling flexibility as well as the nuances of structure learning becomes a natural pedagogical start.
There is a degree of flexibility in the choice of statistical model for defining the transition function. We will explore three options - the noisy or model, the linear logit model, and the complete dependency table model.
@@
$X_{i}(t+1)=\beta_{0}+\sum\limits_{j\in dpa_{d}(i)}\beta^{d}_{j}X_{j}(t)+\sum\limits_{j\in dpa_{s}(i)}\beta^{s}_{j}X_{j}(t+1)+\sum\limits_{\tau\in dpa_{\tau}(i)}\beta^{a}_{\tau}X_{i}(t-\tau)+\sum\limits_{j\in dpa_{z}(i)}\beta^{z}_{j}Z_{j}$ (6)
@@
$p(X_{i}(t+1)=1)=\sigma\left(X_{i}(t+1)\right)$ (7)
with $X_{i}(t+1)$ given by the expression above.
@@
$\begin{array}[]{l}X_{1}(t+1)=\beta_{0}+\beta_{1}Z_{2}+\beta_{2}X_{3}(t+1)+\beta_{3}X_{3}(t+1)Z_{2}+\beta_{4}(1-X_{3}(t+1)Z_{2})+\beta_{5}X_{2}(t)\\ \qquad\qquad\qquad+\beta_{6}(1-X_{2}(t)Z_{2})+\beta_{7}(1-X_{2}(t)X_{3}(t+1))+\beta_{8}X_{2}(t)X_{3}(t+1)Z_{2}\end{array}$
@@
$\begin{array}[]{l}X_{i}(t+1)=\prod\limits_{j\in dpa_{d}(i)}\prod\limits_{k\in dpa_{s}(i)}\prod\limits_{\tau=1,...,p}\prod\limits_{l\in dpa_{z}(i)}\sum_{\alpha\in\left(\mathbb{Z}^{+}_{2}\right)^{4}}\beta_{i,j,k,l,\tau}^{\alpha}\left(W_{jkil}\right)\\ W_{jkil}=\Pi_{jkil}\left(X_{j}(t)X_{k}(t+1)X_{i}(t-\tau)Z_{j}\right)^{\alpha}\\ \qquad\qquad+\left(1-\Pi_{jkil}\right)\left(1-\left(X_{j}(t)X_{k}(t+1)X_{i}(t-\tau)Z_{j}\right)^{\alpha}\right),\,\Pi_{jkil}\in\{0,1\}\end{array}$ (8)
@@
On the other hand, this structure of statistical model presents two structural conditions denoted as local parameter independence and unrestricted multinomial distribution [spiegelhalter1990sequential]. These ensure that for every configuration, that is, every possible combinations of values of the parents of a node, there is an independent parameter vector. This leads to a corresponding combinatorial explosion of parameter vectors in the statistical model. On the other hand, however, the analytical calculation of parameter likelihoods and posterior distributions become possible, facilitating more straightforward evaluation of scoring metrics quantifying the information quality of an entire (D)BN. That is, the marginal likelihood of a structure can be computed without first computing the likelihood of the weights.
@@
$\sum_{j\in dpa(i)}\beta_{j}V_{j}(t+1)=\sum\limits_{j\in dpa_{d}(i)}\beta^{d}_{j}X_{j}(t)+\sum\limits_{j\in dpa_{s}(i)}\beta^{s}_{j}X_{j}(t+1)+\sum\limits_{\tau\in dpa_{\tau}(i)}\beta^{a}_{\tau}X_{i}(t-\tau)+\sum\limits_{j\in dpa_{z}(i)}\beta^{z}_{j}Z_{j}$
@@
$p(X_{i}(t+1)=u^{l}|V_{dpa(i)}(t+1),\theta_{i}):=\theta_{i}^{\xi(V_{dpa(i)}(t+1))},\,\xi\in\Xi_{i},\,\Xi_{i}=\prod_{j\in dpa_{t}(i)}\mathbb{Z}_{m}^{+}\times\prod_{j\in dpa_{z}(i)}\mathbb{Z}_{q}^{+}$ (10)
That is, there is an multi-index that enumerates the entries of $\Xi_{i}$ for each transition $i$. We can see that this presents a highly parametrized model, which will imply significant parametric uncertainty when there are finite data samples. On the other hand, with this highly precise model, the maximum likelihood becomes much more straightforward to compute, as well as the Bayesian scores. Indeed, this is exactly what local parameter independence facilitates – you can compute the likelihood by counting the instances of each transition and dividing by the count of each predecessor configuration. On the other hand, when the total count of every possible predecessor configuration is low, due to unfavorable sample complexity, this can be a rather noisy source of actual validity of that dependence. On the other hand, by letting these parameters take on distributions, in a Bayesian setting, the computation of a posterior for a structure becomes easier, and the uncertainty is available by sampling the posterior, anyway.
@@ 2.1.2 Continuous Variables
Indeed in [koller2009probabilistic, Theorem 7.3-7.4] it is shown that there is a bidirectional equivalence between such a normal joint distribution and normal transition function.
@@
Linear Structural Equation Models (LSEMs) are the most common non-Gaussian DBN for modeling continuous variables. With LSEMs (see, e.g. [bowen2011structural] for a general reference and [peters2014causal] for application to causal inference) presume a general linear structure that is associated with a discretization of a dynamical system:
@@
An SEM could refer to a purely time-instant (static) model, with dependencies $dpa_{s}$ and $dpa_{z}$ only, as in [pmlr-v180-deleu22a, manzour2021integer]. More recently, DBNs more broadly have become interchangeable with SEMs, for instance, the representation in [pamfil2020dynotears] has all dependencies as described here except, it can be argued for simplicity, $Z$.
@@
Broadly speaking, there are a number of popular parametric forms of nonlinear models that can be used from time series literature, e.g. [douc2014nonlinear]. Neural networks have enabled computationally intensive empirical unsupervised time series models [franceschi2019unsupervised, cai2024msgnet]. Simple nonlinear functional models in the SEM statistical community have also been studied [lee2000statistical]. The work [sharma2022scalable] uses splines to model the nonlinear relationships in the transition distributions. The work [kim2004dynamic] uses a kernel nonparametric regression model to learn DBNs for gene regulatory networks.
@@ 3 Learning From Multiple Trajectories
and we are interested in fitting a DBN model to this data. This amounts to defining the specific form of $f$ in (1) and as far as three tasks, as defined in the introduction. More specifically, it amounts to
  1. 1.
    identifying the parents of each $X_{j}(t)$ in the graph $\bar{\mathcal{G}}$, as well as specific functional form of the transition function $f$, of a ground truth
  2. 2.
    Performing inference on the model accurately. For DBNs a common use is for forecasting, but generalization and test error is another common metric.
@@ 3.1 Maximum Likelihood Calculations
In reviewing the literature on learning DBNs from data, it is typical to disregard the distinction of the trajectory sample $\mathcal{T}^{i}$ and the time transition samples $\{X^{(i)}(t),X^{(i)}(t+1)\}$. As far as understanding the meta-methodological cause of this, it appears that this can be said to be due to DBNs being considered not uniquely, but as either a special kind of Bayesian Network, or as splices of the same time series trajectory.
Consider the two methodological components surrounding the study of DBNs, time series analysis and PGMs. For the latter, consider two popular works that are effectively extensions of methods developed for BNs extended to DBNs, the continuous reformulation of the problem into one with adjacency matrices as decision variables, called “NOTEARS” in the static case [zheng2018dags] and “dynotears” [pamfil2020dynotears], as well as the use of “Generative Flow Networks”, a Reinforcement Learning-motivated sampler, for the static case in [pmlr-v180-deleu22a] and the dynamic case in [atanackovic2024dyngfn]. It can be seen that in all of these cases, the likelihood is expressed as,
@@
And similarly, in consulting standard texts on time series analysis with detailed derivations of Likelihood computation for various models, e.g. [douc2014nonlinear, mcquarrie1998regression], we see that in the derivations of the likelihood, the data is considered to be a sequence of observations, that is, a sequence of observations from a stochastic process $\{\hat{X}(0),\hat{X}(1),\hat{X}(2),...,\hat{X}(T)\}$, rather than the general form given in (13), and is fit to (14), just with a simpler expression in the sum index. However, arithmetically, the calculations commute, that is, they are equivalent to the Bayesian Network paradigm. Being equivalent expressions for the likelihood, brought from different perspectives, we can see that the conditional nature of each sample makes the concern as to the appropriate application of the methods to be less as to whether they can adequately computed with sound statistical assumptions, but care as to what one can learn, i.e., learning conditional dependencies rather than long term statistical properties.
3.2 Background on Causal Learning
Conditional Independence and d(irected)-separation
@@ 3.2 Background on Causal Learning
Independence of various sets of variables can be determined by examining d-separation (d means directional) criterion of the (D)BN DAG [ProbabilisticCowell].
@@
Causal Sufficiency
We present the main set of definitions used in the literature from [???, beckers_causal_2021].
A signature $S$ is defined as a tuple $(U,V,R)$, where $U$ is a set of exogenous variables, $V$ is a set of endogenous variables, and $R$ is a function that maps each variable $Y$ in $U\cup V$ to a nonempty set $R(Y)$ of possible values, representing the range of values $Y$ can take. For a vector of variables $\vec{X}=(X_{1},\dots,X_{n})$, $R(\vec{X})$ represents the Cartesian product $R(X_{1})\times\dots\times R(X_{n})$.
A causal model $M$ consists of a pair $(S,F)$, where $S$ represents the signature and $F$ specifies a set of functions. Each function in $F$ associates an endogenous variable $X$ with a structural equation $F_{X}$, which determines the value of $X$ based on the values of other endogenous and exogenous variables. Formally, the function $F_{X}$ maps $R(U\cup V-\{X\})$ to $R(X)$, meaning that $F_{X}$ provides the value of $X$ given the values of all other variables in the set $U\cup V$.
@@
Causal discovery and Inference
@@
3.3 Considerations from Axioms of Causal Learning
DBNs, compared to BNs, contain both time-varying as well as static variables. While the in-time transition structure is still a DAG, suggesting that formally many of the same principles regarding inference as well as structure and weight learning in BNs carry over to DBNs, the presence of time, especially when long trajectories are expected, adds significant complications.
This indicates the utility of including both static $Z$ and dynamic $X$ variables in the model for elucidation. Instead one can consider a new trajectory as a new sample of $\hat{Z}$, which itself samples $X(0)\sim\pi(X(0)|Z)$ then, $X(0),X(1),...,X(T)$. As such, one has $T$ samples in order to learn $P\left(X(t+1)|X(t),\hat{Z}\right)$. However, what can be said about how informative a marginal trajectory is towards learning $P\left(X(t+1)|X(t)\right)$, that is, the marginal conditional over the population of $Z$?
As far the required observations for causal identification, this is that there exists at least one $Z$ such that for all trajectories $Z$ is observed, and $Z$ is in the parent of some $X(t)$. We can consider that the classic Randomized Clinical Trial is exactly that $Z_{H}\in\{0,A\}$ and then testing for $p(X(t+1)|X(t),Z_{\setminus},Z_{H}=0)\neq p(X(t+1)|X(t),Z_{\setminus},Z_{H}=A)$, with a null and alternative hypothesis and $Z_{\setminus}$ as other covariates, assumed to me completely independent of $Z_{H}$.
@@ 3.4 Closed System Graph Causal Identification Model and Likelihood Information
wherein $\epsilon$ is a zero mean error term. This additive noise model formulation has been leveraged to establish results on the identifiability of the structure $\bar{\mathcal{G}}$ [peters2014causal, hoyer2008nonlinear].
@@
Finally, let us investigate in more detail the information regarding the DBN dynamics given by (14).
@@
from which we can see the intuition of the circumstance. Asymptotically, the second term approaches the stationary distribution, and the independence assumption becomes valid. Indeed, the structure of the DAG and overall model is expected to be such that the system is long run ergodic. Otherwise, we can consider that for $T$ much longer than the mixing time, this assumption is also valid for most of the transitions. For shorter times of trajectories we can see that:
@@
    where $V_{i}$ could be any predecessor in the graph for $X_{j}(t+1)$.
@@
We can finally conclude from this analysis that the distinction in time transitions and trajectories is only relevant as far as which conditional dependencies they present. Fundamentally, the overall population prior observed for the conditioned variable determines the usefulness of the observations for learning statistical long run averages of the process.